OpenStudy (anonymous):

A tuition scale for a state university shows an increase of $1,100 each year. If the initial cost was $27,000. What will the cost of tuition be in twenty two years? In complete sentences, explain how an arithmetic sequence was applied?

OpenStudy (anonymous):

ok so in this case, our first term a1 is 27,000. Now every year, we add by a constant 1,100. So the numbers which are the cost of a state university each year gives us an arithmetic sequence since then we'll have a list of numbers which are increasing by 1,100 every term. since the first term is 27000 and the difference is 1100 then the nth term of our sequence(or simply the cost of tuition in the nth year) is given by a_n=27000+(n-1)1100, so the cost in 20 years is a_20=27000+(19)1100=27000+20900=47900 dollars is the answer

OpenStudy (anonymous):

thank u !!!!